OpenStudy (anonymous):

What is the solution of this equation? WORK SHOWN! t + 16= 7

OpenStudy (cicilybailey14):

What plus 16 equals 7?

OpenStudy (anonymous):

Read the problem.

OpenStudy (texaschic101):

try subtracting 16 from both sides

OpenStudy (anonymous):

I need the work shown

OpenStudy (reni_love_ela):

no im trippin 23

OpenStudy (cicilybailey14):

@Reni_Love_ELA she said work shown. I was going to show work -_-

OpenStudy (anonymous):

is that correct

OpenStudy (cicilybailey14):

Let's solve your equation step-by-step. t+16=7 Step 1: Subtract 16 from both sides. t+16−16=7−16 t=−9 Answer: t=−9

OpenStudy (reni_love_ela):

omg i keep reading this worng

OpenStudy (anonymous):

-9 + 16 = 7

OpenStudy (cicilybailey14):

T=-9

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

t + 16 = 7 t = 7 - 16 t = - 9

OpenStudy (cicilybailey14):

welcome!